Circular No. 8183

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2003 (130) 1 W. J. Merline, Southwest Research Institute (SwRI); P. M. Tamblyn,

Binary Astronomy, Golden, CO, and SwRI; C. Dumas, Jet Propulsion Laboratory; L. M. Close, University of Arizona; C. R. Chapman, SwRI; and F. Menard, Observatoire de Grenoble, report the discovery on Aug. 15.6 UT of a satellite of minor planet (130) Elektra from K′- and H-band imaging with the 10-m Keck II Telescope (+ NIRC2/AO adaptive optics system) on Mauna Kea. On Aug. 15.6417, the satellite was at separation 0′′.72 (projected separation 1170 km) and position angle 70^{o}. The satellite was confirmed by the same team, also at Keck II, on Aug. 17.6. The K′-band brightness difference is ∼ 8.5 mag, giving an estimated diameter of the satellite of ∼ 4 km.

SUPERNOVA 2003hd IN MCG –04-5-10 M. Hamuy, N. Morrell, and J. Thomas-Osip, Carnegie Observatories,

report that a spectrum (range 360–900 nm) of SN 2003hd (cf. IAUC 8179), obtained on Aug. 15.28 UT with the Las Campanas Clay 6.5-m telescope (+ LDSS-2), shows it to be a young type-II supernova with a blue continuum having incipient Hα and Hβ broad P-Cyg lines.

COMET C/2003 E1 (NEAT)

Orbital elements from MPC 49276:

Epoch = 2004 Feb. 5.0 TT

T = 2004 Feb. 13.6332 TT  ω = 103.^{o}8634
e = 0.763499  Ω = 137.0688 }2000.0
q = 3.245045 AU  i = 33.5375

a = 13.721077 AU n^{o} = 0.0193920 P = 50.83 years

COMET C/2002 Y1 (JUELS-HOLVORCEM) Total visual magnitude estimates: July 3.74 UT, 9.0 (D. A. J. Seargent,

The Entrance, N.S.W., 25×100 binoculars); 27.68, 10.6 (M. Mattiazzo, Wallaroo, South Australia, 25×100 binoculars).

CORRIGENDUM On IAUC 8182, ‘Comet C/2003 H1 (LINEAR)’, line 1, for parabolic

read nearly-parabolic

2003 August 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green